In which gas there is are same number of molecules as present in 16 gm oxygen- (1)16 g O3(2)16 g SO2(3)32 gm SO2(4)all of these Explain????

Grade:11

1 Answers

Vikas TU
7 years ago
No. of molecules in Oxygen, => 16/32 = Na/2
No. of molecules in Ozone,=> 16/48 = Na/3
No. of molecules in SO2,=> 16/64 => Na/4
No. of molecules in 32 gm SO2, => 32/64 => Na/2